OpenStudy (anonymous):

second derivative of y3+y-2 x2 y=12

OpenStudy (yttrium):

\[y^3+y-2x^2y=12\] Do you need to find y''?

OpenStudy (anonymous):

with respective to x & y both

OpenStudy (yttrium):

If so, \[3y^2y'+y'-2(x^2y'+2xy')=0\]\[3y^2y'+y'-2x^2y'-4xy=0\] Hence. \[y'=\frac{ 4xy }{ 3y^2-2x^2+1 }\] Second derivative. \[y''\frac{ (3y^2-2x^2+1)[4(xy'+y)]-(4xy)[6yy'-4x] }{ (3y^2-2x^2+1)^2 }\]

OpenStudy (anonymous):

thank q